### v1

@@ Abstract
3D human pose estimation aims to reconstruct the human skeleton of all the individuals in a scene by detecting several body joints. The creation of accurate and efficient methods is required for several real-world applications including animation, human-robot interaction, surveillance systems or sports, among many others. However, several obstacles such as occlusions, random camera perspectives, or the scarcity of 3D labelled data, have been hampering the models’ performance and limiting their deployment in real-world scenarios. The higher availability of cameras has led researchers to explore multi-view solutions due to the advantage of being able to exploit different perspectives to reconstruct the pose.
Thus, the goal of this survey is to present an overview of the methodologies used to estimate the 3D pose in multi-view settings, understand what were the strategies found to address the various challenges and also, identify their limitations. Based on the reviewed articles, it was possible to find that no method is yet capable of solving all the challenges associated with the reconstruction of the 3D pose. Due to the existing trade-off between complexity and performance, the best method depends on the application scenario. Therefore, further research is still required to develop an approach capable of quickly inferring a highly accurate 3D pose with bearable computation cost. To this goal, techniques such as active learning, methods that learn with a low level of supervision, the incorporation of temporal consistency, view selection, estimation of depth information and multi-modal approaches might be interesting strategies to keep in mind when developing a new methodology to solve this task.
keywords:
[…]
@@ 1.1 Previous literature reviews
The increased availability of multi-camera setups has prompted further research on the benefits of having various camera perspectives. As a result, several surveys, besides analysing single-view methods, also cover multi-view approaches [96, 55, 22, 70, 85, 62, 8, 100, 49, 47]. The literature review of Holte et al. [34] is the only one found, solely dedicated to multi-view methodologies. According to the analysed surveys, most of the problems observed in monocular settings are the same for multi-view, namely, the lack of annotated in-the-wild datasets or with rare poses [55, 70, 8, 22, 49, 85, 100, 62], the difficulty in identifying the several poses in crowded scenarios, and the occlusions [96, 47, 100, 49, 34]. Even though, overcoming the occlusions problem benefits from the different view perspectives, it still remains a challenge in some situations. Besides, the increase in the number of viewpoints leads to the need for more complex models to be able to deal with the information acquired from all perspectives. This translates into slower inference runtime, forcing a balance between complexity and speed [8, 47, 100, 49]. Furthermore, the bad image quality in terms of focus or blur [100, 47, 49, 34], the inaccuracies of 2D pose estimations which affect the models that use 2D-3D lifting [49, 34] are also, factors negatively influencing the quality of the predictions. So, there are still a lot of opportunities for improvement, such as considering the human interactions either with other humans or the surrounding environment [70, 85, 100, 34], the increase of the generality in order for the same method to be suitable across multiple applications and independent of the presented viewpoints [47, 49, 34]. Also, several surveys recommend the use of neural architecture search for a better and more efficient design of neural networks [85, 47, 100, 49], taking advantage of temporal information, even the adaptation of methods from monocular to multi-view [47, 22], the use of transfer learning [55] and data augmentation [100]. Finally, to combat the scarcity of annotated datasets, several strategies are suggested like using unsupervised, semi-supervised or weakly-supervised models [55, 85], or the use of active learning to alleviate the human workload [62].
@@ 1.2 Search Process
The rest of this literature review is organised in the following way: Chapter 2 - Datasets, Chapter 3 - Evaluation Metrics, Chapter 4 - Multi-view approaches, Chapter 5 - Multi-modal approaches, Chapter 6 - Conclusion.
@@ 2 Datasets
Datasets Year Size Nº cameras Nº subjects Characteristics
@@ 3 Evaluation Metrics
$º\text{MPJAE}=\frac{1}{N}\sum_{i=1}^{N}\left|\left(x_{i}-x_{i}^{{}^{\prime}}\right)\mod\pm 180\text{\textordmasculine}\right|$ (6)
@@
[…]
@@ 4.1 Single-person approaches
In this section, it’s presented methods developed to retrieve the 3D pose of scenarios with just one individual.
Sigal et al. [74] introduced the HumanEva datasets to combat the lack of datasets for multi-view 3D pose estimation and provide a benchmark so that developed methods could be compared fairly. Thus, to provide intuition on how to work with their dataset, a baseline model was proposed which consisted of a Bayesian filtering method whose parameters were optimised using Sequential Importance Resampling and Annealed Particle Filtering.
@@
As for Remelli et al. [65], they presented an efficient method of reconstructing the pose without substantially increasing the computational cost. To do so, the authors introduced a faster version of the Direct Linear Transform method to lift the predicted 2D skeletons to 3D poses. The 2D poses are generated by exploring 3D geometry and camera projections.
@@
The computational load imposed by Convolutional Neural Network (CNN)-based models may limit their application in real-world scenarios. To surpass this problem, Hwang et al. [36] designed a system consisting of several edge devices and a central server. In which each edge device was responsible for estimating the 2D pose of the received image using a CNN and then, sending it along with the respective timestamp to the central server. The central server, in turn, triangulates using Direct Linear Transform (DLT) the received 2D poses that were detected at roughly the same time to produce the 3D pose. Hence, this system distributes the computational burden of the CNN among edge devices and also reduces data traffic by transmitting only the 2D poses and timestamps rather than RGB images, making it more suitable for real-time applications in the real-world.
4.1.1 Methods under different supervision levels
The scarcity of labelled datasets significantly restricts the models’ outcomes. Therefore, unsupervised, weakly-supervised, semi-supervised and self-supervised methods have been proposed to combat the dependence on 3D labelled datasets.
@@ 4.1.1 Methods under different supervision levels
In the research work carried out by Zhao et al. [99], a new loss function, Triangulation Residual Loss, has been proposed, whose objective is to minimise the total distances between view rays and 3D pose estimation retrieved through triangulation. This novel loss function aims to enforce multi-view geometric consistency, facilitating the efficient self-supervised training of the model. As for Jenni and Favaro [40], they proposed a pre-training technique using self-supervised learning, to determine the synchronisation between two images and whether one image is inverted horizontally in relation to the other, with the objective of learning a meaningful representation suitable to be used by a network to estimate the 3D pose. Also, the authors suggest removing the background for static images to prevent irrelevant features from being learnt during the self-supervised learning task. In addition, Liu et al. [53] also, introduced a method which relies only on two views to reconstruct the pose, without requiring to know any camera configurations. So, their method after extracting the 2D poses of each image, used a self-supervised 3D regression network able to produce virtual views using orthogonal projections of the human body, allowing the model to thoroughly understand the human body spatial structure and the transformations necessary to project one perspective onto another. Concerning the work of Yin et al. [95], a self-supervised method incorporating temporal convolution blocks and spatio-temporal attention mechanisms has been investigated. The methodology employed involved the exploration of re-projection and view consistency methods, which enabled an accurate retrieval of the 3D pose of infants without the need for 3D labels or definition of camera calibration parameters.
@@
Papers Year MPJPE (mm) $\downarrow$
Rhodin et al. [67] 2018 131.70
Tang et al. [79] 2018 99.70
@@ 4.2 Multi-person approaches
This section presents the solutions developed to determine 3D poses in multi-person scenarios.
@@ 4.2.1 Geometric constraint-based methods
In [6], the authors based their work on their previous work [5]. Nonetheless, they used different body parametrisations by exploring the use of a 3DPS with several potential functions and also, to learn the model’s parameters was used a structured Support Vector Machine (SVM) to adequately weight the different potential functions. Zhu et al. [102] also, explored the PS model, but for human motion tracking. To accomplish that, the researchers first, subtracted the background, then, used Flexible Mixtures of Parts (FMP), which is based on PS, for foreground learning to detect the human body parts and finally, to track the body parts, it was utilised Annealed Particle Filter (APF). The latest work by Belagiannis et al. [6], as mentioned, relied on a structured SVM to optimise the model parameters. In contrast, Schwarcz and Pollard [71], to simplify the hyperparameters optimisation and improve the quality of the poses estimated with triangulation, used, to determine the 3D pose, a Conditional Random Field (CRF) as a factor graph, the 3D limb locations as variables, and the limb position priors, collision terms and temporal smoothing on those joints as factors. Furthermore, for determining the 2D poses for each view, Schwarcz and Pollard took advantage of the OpenPose library [11]. On the other hand, Ershadi et al. [27] employed a bottom-up approach starting by leveraging from the DeeperCut model [38] to detect the 2D body parts. Next, a 3D search space is constructed with all detected joints and a Gaussian Mixture Model (GMM) is applied to cluster the points in order to obtain the number of people in the scene. To infer the 3D pose, a fully connected pairwise CRF with its pairwise terms defined in the 3D space and a loopy belief propagation are utilized. To alleviate the computational cost of deep neural networks developed to estimate the 3D pose, Wang et al. [84] proposed LHPE-nets. The LHPE-nets incorporate a low-span network to increase the speed of the start of the training process for the 2D pose prediction as well as a residual deep neural network trained on low-resolution data which showed to enhance network scalability and performance when compared to a ResNet-34.
@@
To combat the inefficiency of PS-based models, Dong et al. [24] proposed a multi-way matching algorithm which leverages epipolar geometry, appearance similarity and cycle consistency to reduce the state space and eliminate false detections, resulting in a more efficient matching of 2D poses between views. Later, in [23], the researchers extend their approach to also, track the poses by introducing temporal tracking and Riemannian Extended Kalman filtering. However, their methods present some limitations, like a 2D pose can be deemed an outlier if it only appears in one view, or if there are fast movements, the tracking algorithm is unreliable. Thus, their method could not be used in applications requiring quick movement monitoring, such as sports tracking. Bridgeman et al. [9], to improve the speed of the 2D pose association across views, employed a fast greedy algorithm based only on geometric relations and then, used triangulation to generate the 3D skeletons. Although, the presented method is faster than PS-based methods, in scenarios with a small number of cameras the use of PS for joint estimation has been proven to be more precise than triangulation [23, 9]. Demonstrating the existing trade-off between speed and accuracy [23].
@@ 4.2.3 Plane sweep stereo-based models
Table 4: Table summarising the results reported in the revised publications for the Campus [5] and Shelf [5] datasets
Papers Year Campus — PCP (%) $\uparrow$ Shelf — PCP (%) $\uparrow$

### v2

@@ Abstract
3D human pose estimation involves reconstructing the human skeleton by detecting the body joints. Accurate and efficient solutions are required for several real-world applications including animation, human-robot interaction, surveillance, and sports. However, challenges such as occlusions, 2D pose mismatches, random camera perspectives, and limited 3D labelled data have been hampering the models’ performance and limiting their deployment in real-world scenarios. The higher availability of cameras has led researchers to explore multi-view solutions to take advantage of the different perspectives to reconstruct the pose.
Most existing reviews have mainly focused on monocular 3D human pose estimation, so a comprehensive survey on multi-view approaches has been missing since 2012. According to the reviewed articles, the majority of the existing methods are fully-supervised approaches based on geometric constraints, which are often limited by 2D pose mismatches. To mitigate this, researchers have proposed incorporating temporal consistency or depth information. Alternatively, working directly with 3D features has been shown to completely overcome this issue, albeit at the cost of increased computational complexity. Additionally, models with lower levels of supervision have been identified to help address challenges such as annotated data scarcity and generalisation to new setups. Therefore, no method currently addresses all challenges associated with 3D pose reconstruction, and a trade-off between complexity and performance exists. Further research is needed to develop approaches capable of quickly inferring a highly accurate 3D pose with bearable computation cost. Techniques such as active learning, low-supervision methods, temporal consistency, view selection, depth information estimation, and multi-modal approaches are strategies to keep in mind when developing a new methodology to solve this task.
Keywords:3D Human Pose Estimation  $\cdot$ Multi-view  $\cdot$ Supervision level  $\cdot$ Temporal consistency  $\cdot$ Multi-modal
@@ 1.1 Previous literature reviews
The increased availability of multi-camera setups has prompted further research on the benefits of having various camera perspectives. As a result, several surveys, besides analysing single-view methods, also cover multi-view approaches [96, 54, 23, 69, 85, 61, 7, 100, 48, 46]. The literature review of Holte et al. [33] is the only one found, solely dedicated to multi-view methodologies. According to the analysed surveys, most of the problems observed in monocular settings are the same for multi-view, namely, the lack of annotated in-the-wild datasets or with rare poses [54, 69, 7, 23, 48, 85, 100, 61], the difficulty in identifying the several poses in crowded scenarios, and the occlusions [96, 46, 100, 48, 33]. Even though, overcoming the occlusions problem benefits from the different view perspectives, it still remains a challenge in some situations. Besides, the increase in the number of viewpoints leads to the need for more complex models to be able to deal with the information acquired from all perspectives. This translates into slower inference runtime, forcing a balance between complexity and speed [7, 46, 100, 48]. Furthermore, the bad image quality in terms of focus or blur [100, 46, 48, 33], the inaccuracies of 2D pose estimations which affect the models that use 2D-3D lifting [48, 33] are also, factors negatively influencing the quality of the predictions. So, there are still a lot of opportunities for improvement, such as considering the human interactions either with other humans or the surrounding environment [69, 85, 100, 33], the increase of the generality in order for the same method to be suitable across multiple applications and independent of the presented viewpoints [46, 48, 33]. Also, several surveys recommend the use of neural architecture search for a better and more efficient design of neural networks [85, 46, 100, 48], taking advantage of temporal information, even the adaptation of methods from monocular to multi-view [46, 23], the use of transfer learning [54] and data augmentation [100]. Finally, to overcome the scarcity of annotated datasets, several strategies are suggested like using unsupervised, semi-supervised or weakly-supervised models [54, 85], or the use of active learning to alleviate the human workload [61].
@@ 1.2 Search Process
The rest of this literature review is organised in the following way: Chapter 2 - Datasets, Chapter 3 - Evaluation Metrics, Chapter 4 - Multi-view approaches, Chapter 5 - Multi-modal approaches, Chapter 6 - Discussion, Chapter 7 - Conclusion.
@@ 2 Datasets
Datasets Year
Size
(Nº frames)
Nº cameras Nº subjects Characteristics
@@ 3 Evaluation Metrics
$\text{MPJAE}=\frac{1}{N}\sum_{i=1}^{N}\left|\left(x_{i}-x_{i}^{{}^{\prime}}\right)\mod\pm 180\text{º}\right|$ (6)
@@
$\mod$ is the modulus operator, the term $\text{mod}\pm 180$º applies to angles brings them into the range of $[-180\text{º},+180\text{º}]$.
@@ 4 Multi-view approaches
The proposed taxonomy, see Figure 3, is as follows: single-person approaches were divided into fully-supervised, and methods from semi-supervised to unsupervised. The majority of the existing approaches use geometric relationships to infer the 3D pose, nonetheless, some researchers have opted for fully-supervised methods, as others have chosen methods with lower levels of supervision to enhance the generalisation capabilities of their methods and possibly improve their models’ performances. Regarding multi-person approaches, those were categorised taking into consideration the type of approach to use to infer the 3D pose from the images. Thus, the methods were categorised into geometric constraint-based if their implementation was based on geometric relations between the camera perspectives or the human body structure. Then, the methodologies under the category of voxel-based methods are approaches which don’t rely on 2D pose estimates to determine the 3D pose and use the voxel features to infer the 3D pose. Finally, plane sweep stereo-based methods are approaches that have used the plane sweep stereo network to incorporate depth information into the estimation of the 2D poses to minimise the impact of mismatches.
Figure 3: Proposed taxonomy for multi-view 3D pose estimation
@@ 4.1 Single-person approaches
In this section, it’s presented methods developed to retrieve the 3D pose of scenarios with just one individual. These methods are divided into two categories: Fully-supervised methods (Section 4.1.1), and From Semi-supervi-sed to Unsupervised methods (Section 4.1.2) which include semi-supervised, weakly-supervised, self-supervised and unsupervised methods. This classification is based on the quantity of labelled data used during training. Therefore, models utilizing fully annotated datasets in their training process are designated as fully-supervised methods, while those utilising little to no annotated datasets are designated under the second category. The difference between semi-supervised, weakly-supervised, self-supervised and unsupervised methods is mainly due to the amount of labelled data used for model training. Therefore, semi-supervised methods are trained with more labelled data than weakly-supervised, which in turn are trained with more annotated data than unsupervised methods, which are trained without any labelled data. Self-supervised approaches are a sub-group of unsupervised methods in which implicit labels are generated to help understand the unstructured data. Unsupervised methods, generally, learn to analyse and understand the data structure without relying on labelled data, clustering it into different groups based on the observed characteristics and relationships between the different data examples.
4.1.1 Fully-supervised methods
Sigal et al. [73] introduced the HumanEva datasets to address the lack of datasets for multi-view 3D pose estimation and provide a benchmark so that developed methods could be compared fairly. Thus, to provide intuition on how to work with their dataset, a baseline model was proposed which consisted of a Bayesian filtering method whose parameters were optimised using Sequential Importance Resampling and Annealed Particle Filter (APF).
@@ 4.1.1 Fully-supervised methods
As for Remelli et al. [64], they presented an efficient method of reconstructing the pose without substantially increasing the computational cost. To do so, the authors introduced a faster version of the Direct Linear Transform (DLT) method to lift the predicted 2D skeletons to 3D poses. The 2D poses are generated by exploring 3D geometry and camera projections.
@@
The computational load imposed by Convolutional Neural Network (CNN)-based models may limit their application in real-world scenarios. To surpass this problem, Hwang et al. [35] designed a system consisting of several edge devices and a central server. In which each edge device was responsible for estimating the 2D pose of the received image using a CNN and then, sending it along with the respective timestamp to the central server. The central server, in turn, triangulates using DLT the received 2D poses that were detected at roughly the same time to produce the 3D pose. Hence, this system distributes the computational burden of the CNN among edge devices and also reduces data traffic by transmitting only the 2D poses and timestamps rather than RGB images, making it more suitable for real-time applications in the real-world.
4.1.2 From Semi-supervised to Unsupervised methods
The scarcity of labelled datasets significantly restricts the models’ outcomes. Therefore, unsupervised, weakly-supervised, semi-supervised and self-supervised methods have been proposed to surpass the dependence on 3D labelled datasets.
@@ 4.1.2 From Semi-supervised to Unsupervised methods
In the research work carried out by Zhao et al. [99], a new loss function, Triangulation Residual Loss, has been proposed, whose objective is to minimise the total distances between view rays and 3D pose estimation retrieved through triangulation. This novel loss function aims to enforce multi-view geometric consistency, facilitating the efficient self-supervised training of the model. As for Jenni and Favaro [39], they proposed a pre-training technique using self-supervised learning, to determine the synchronisation between two images and whether one image is inverted horizontally in relation to the other, with the objective of learning a meaningful representation suitable to be used by a network to estimate the 3D pose without relying on knowing any camera parameter. Also, the authors suggest removing the background for static images to prevent irrelevant features from being learnt during the self-supervised learning task. In addition, Liu et al. [52] also, introduced a method which relies only on two views to reconstruct the pose, without requiring to know any camera configurations. So, their method after extracting the 2D poses of each image, used a self-supervised 3D regression network able to produce virtual views using orthogonal projections of the human body, allowing the model to thoroughly understand the human body spatial structure and the transformations necessary to project one perspective onto another. Concerning the work of Yin et al. [95], a self-supervised method incorporating temporal convolution blocks and spatio-temporal attention mechanisms has been investigated. The methodology employed involved the exploration of re-projection and view consistency methods, which enabled an accurate retrieval of the 3D pose of infants without the need for 3D labels or definition of camera calibration parameters.
@@
Type of method Paper Year MPJPE (mm) $\downarrow$
Fully-supervised Tang et al. [78] 2018 99.70
@@
Jiang et al. [41] 2023 27.80
Wan et al. [82] 2023 21.10
Cai et al. [9] 2024 15.10
Cai et al. [9] using the 2D ground-truth poses 2024 7.90
Semi-supervised Rhodin et al. [66] 2018 131.70
Self-supervised Jenni and Favaro [39] 2021 64.90
Liu et al. [52] 2021 22.50
@@ 4.2 Multi-person approaches
This section presents the solutions developed to determine 3D poses in multi-person scenarios. As such, this section will be divided according to the proposed taxonomy, into geometric constraint-based methods (Section 4.2.1), voxel-based methods (Section 4.2.2) and plane sweep stereo-based methods (Section 4.2.3).
The approaches have different characteristics, thus, geometric constraint-based methods leverage from using constraints regarding the structure of the human body, epipolar geometry, geometric relationships between views, and similar conditions. Plane sweep stereo-based methods also use geometric constraints and try to relate 2D poses from various perspectives to create the final 3D pose, nonetheless, they rely on depth information calculated using the plane sweep stereo-based model. The voxel-based methods have a more distinct approach, in which the 3D pose is regressed based on a set of features obtained in the 3D space and not on the 2D space which is the most common approach by the other two mentioned categories, preventing these models from being subjected to the performance of a 2D pose estimator and matching processes. Nonetheless, usually are more computationally expensive and slower.
@@ 4.2.1 Geometric constraint-based methods
In [4], the authors based their work on their previous work [5]. Nonetheless, they used different body parametrisations by exploring the use of a 3DPS with several potential functions and also, to learn the model’s parameters was used a structured Support Vector Machine (SVM) to adequately weight the different potential functions. Zhu et al. [102] also, explored the PS model, but for human motion tracking. To accomplish that, the researchers first, subtracted the background, then, used Flexible Mixtures of Parts (FMP), which is based on PS, for foreground learning to detect the human body parts and finally, to track the body parts, it was utilised APF. The latest work by Belagiannis et al. [4], as mentioned, relied on a structured SVM to optimise the model parameters. In contrast, Schwarcz and Pollard [70], to simplify the hyperparameters optimisation and improve the quality of the poses estimated with triangulation, used, to determine the 3D pose, a Conditional Random Field (CRF) as a factor graph, the 3D limb locations as variables, and the limb position priors, collision terms and temporal smoothing on those joints as factors. Furthermore, for determining the 2D poses for each view, Schwarcz and Pollard took advantage of the OpenPose library [10]. On the other hand, Ershadi et al. [27] employed a bottom-up approach starting by leveraging from the DeeperCut model [37] to detect the 2D body parts. Next, a 3D search space is constructed with all detected joints and a Gaussian Mixture Model (GMM) is applied to cluster the points in order to obtain the number of people in the scene. To infer the 3D pose, a fully connected pairwise CRF with its pairwise terms defined in the 3D space and a loopy belief propagation are utilized. To alleviate the computational cost of deep neural networks developed to estimate the 3D pose, Wang et al. [84] proposed LHPE-nets. The LHPE-nets incorporate a low-span network to increase the speed of the start of the training process for the 2D pose prediction as well as a residual deep neural network trained on low-resolution data which showed to enhance network scalability and performance when compared to a ResNet-34.
@@
To deal with the inefficiency of PS-based models, Dong et al. [24] proposed a multi-way matching algorithm which leverages epipolar geometry, appearance similarity and cycle consistency to reduce the state space and eliminate false detections, resulting in a more efficient matching of 2D poses between views. Later, in [25], the researchers extend their approach to also, track the poses by introducing temporal tracking and Riemannian Extended Kalman filtering. However, their methods present some limitations, like a 2D pose can be deemed an outlier if it only appears in one view, or if there are fast movements, the tracking algorithm is unreliable. Thus, their method could not be used in applications requiring quick movement monitoring, such as sports tracking. Bridgeman et al. [8], to improve the speed of the 2D pose association across views, employed a fast greedy algorithm based only on geometric relations and then, used triangulation to generate the 3D skeletons. Although, the presented method is faster than PS-based methods, in scenarios with a small number of cameras the use of PS for joint estimation has been proven to be more precise than triangulation [25, 8]. Demonstrating the existing trade-off between speed and accuracy [25].
@@ 4.2.3 Plane sweep stereo-based models
Table 4: Table summarising the results reported in the revised publications for the Campus [5] and Shelf [5] datasets.
Type of method Paper Year Campus — PCP (%) $\uparrow$ Shelf — PCP (%) $\uparrow$
@@
Plane sweep stereo-based Lin and Lee [49] 2021 98.40 93.70 99.00 97.00 99.30 96.50 98.00 97.90
de França Silva et al. [19] 2022 96.90 87.80 88.90 91.20
Zhou et al. [101] 2022 98.60 94.00 99.20 97.30 99.40 97.30 98.20 98.30